OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

if you use \[\frac{y_{1}-y_{2}}{x_{1}-x_{2}}=m=slope\] then apply the slope and one set of coordinates into \[y - y_{1}=m(x - x_{1})\] solve for y then you will have your y = mx +b equation :) hope this helps

OpenStudy (anonymous):

Yes, it helps, thank you very much! ^_^

OpenStudy (anonymous):

no problem :)

OpenStudy (uri):

I like to solve these Questions but it was in you luck tosay @hobbs978 xD

OpenStudy (uri):

today*

OpenStudy (anonymous):

:)